If edging cost $2.32 per 12-inch stone, and you want a double layer of edging around your flower bed that is 6 yards by 1 yard. How much will edging you flower bed cost?


Answer by Abbey(339)

convert the flower bed from yardage to feet by multiplying by 3 (3 feet in a yard)
Length = 6*3 or 18 feet
Width = 1*3 or 3 feet
then use the formula:
Perimeter = 2*length + 2*width
Perimeter = 2(18)+2(3)
perimeter = 36+6
Perimeter = 42 feet
But you want a double layer of edging, so you will need:
2*42 feet of edging
84 feet of edging.
If edging costs 2.32 per foot (12"stone = 1 foot)
84 feet*$2.32=$194.88 total